OpenStudy (amtran_bus):

How many 1 inch square tiles would it take to cover the floor of a closet that has dimensions 5 feet by 4 feet? (1 foot = 12 inches)

OpenStudy (mathmale):

Find the area (in square feet) of the closet floor, which itself is 5 ft. by 4 ft. Next, convert this area to square inches, using the info that 144 sq. in. = 1 sq. ft.

OpenStudy (amtran_bus):

So area is 20 sq ft? So 2880

OpenStudy (amtran_bus):

Thanks.

OpenStudy (sshayer):

5 ft=5*12=?inches 4 ft=4*12=? inches area=? number of tiles=area of floor in sq.inches/area of one tile

OpenStudy (amtran_bus):

Thanks! I got it as 2880

OpenStudy (mathmale):

Nice work!